New Mexico, Oklahoma and Texas are known for growing peanuts. Imagine that a new disease attacks the peanut plants and damages a large part of the crop. What happens to the consumer surplus in the market for peanuts and for peanut butter as a result of this change?

Consumer surplus in the market for peanuts increases, and the consumer surplus in the market for peanut butter increases.

Consumer surplus in the market for peanuts increases, and the consumer surplus in the market for peanut butter decreases

Consumer surplus in the market for peanuts decreases, and the consumer surplus in the market for peanut butter decreases.

Consumer surplus in the market for peanuts decreasses, and the consumer surplus in the market for peanut butter increases.

In her spare time, Kathy babysits her neighbors' kids for extra income. Her neighbors will pay her $15 per hour. Kathy is willing to babysit for the first hour for $5, the second hour for $7, the third hour for $10, the fourth hour for $14, the fifth hour for $18, the sixth hour for $25. Assume that Kathy is rational in deciding how many hours to spend babysitting. Her producer surplus is

New medical research is published and indicates that drinking orange juice improves the functioning of the immune system. Assuming that everything else remains constant, what would happen to the equilibrium price and producer surplus in the market for orange juice?

The price will increase, and producer surplus will decrease.

The price will decrease, and producer surplus will increase.

The price will decrease, and producer surplus will decrease.

The price will increase, and producer surplus will increase.
Answer #1

Consumer surplus in the market for peanuts decreases, and the consumer surplus in the market for peanut butter decreases.
(Supply of both peanut and peanut butter will decrease which will decrease the consumer surplus in both markets.)

As Kathy is rational, so she will babysit for first four hours. Her producer surplus = (15 - 5) + (15 - 7) + (15 - 10) + (15 - 14) = 10 + 8 + 5 + 1 = $24

The price will increase, and producer surplus will increase.
(Due to new medical research, demand for orange juice will increase which will increase the price and quantity sold of orange juice so producer surplus will increase.)